OpenStudy (anonymous):

MEDAL + FAN FOR RIGHT ANSWER!!!! Portia drew a line on the coordinate plane that passed through the points (-3,19) and (2,9). Which equation represents the line that Portia drew?

OpenStudy (dan815):

y=mx+b m=slope b=y-intercept

OpenStudy (dan815):

slope = \[\frac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

A. y = -2x + 13 B. y = -1/2x + 7 C. y = -1/2x + 13 D. y = -2x + 7

OpenStudy (anonymous):

id say b

OpenStudy (dan815):

y-y1=m(x-x1) is a line with slope m that passes through the point (x1,y1) you can rearrange this to be y=mx-mx1+y1 so b=-mx1+y1

OpenStudy (anonymous):

y=-2x+13

OpenStudy (anonymous):

i meant a

OpenStudy (anonymous):

got them confused

OpenStudy (anonymous):

thanks guys